# EnsAI: An Emulator for Atmospheric Chemical Ensembles

### v1

@@ Abstract
Ensemble-based methods for data assimilation and emission inversions are a popular way to encode flow-dependency within the model error covariance. While most ensemble methods do not require the use of an adjoint model, the need to repeatedly run a geophysical model to generate the ensemble can be a significant computational burden. In this paper, we introduce EnsAI, a new AI-based ensemble generation system for atmospheric chemical constituents. When trained on an existing ensemble for ammonia generated by the GEM-MACH air quality model, it was shown that the ensembles produced by EnsAI can accurately reproduce the meteorology-dependent features of the original ensemble, while generating the ensemble 3,300 times faster than the original GEM-MACH ensemble. While EnsAI requires an upfront cost for generating an ensemble used for training, as well as the training itself, the long term computational savings can greatly exceed these initial computational costs. When used in an emissions inversion system, EnsAI produced similar inversion results to those in which the original GEM-MACH ensemble was used while using significantly less computational resources.
@@ 1 Introduction
While having an accurate estimate of a model’s error covariance is of high importance in many applications, actually obtaining such an estimate is often difficult. Earlier numerical weather prediction (NWP) assimilation systems often used static forecast error covariances as this was more computationally feasible at the time. But as more computational resources became available, more weather centers were able to incorporate flow-dependent error covariances within their systems. One popular method for doing this was by using four-dimensional variational (4DVar) (Thepaut and Courtier, 1991; Zupanski, 1993; Rabier et al., 2000). However, 4DVar’s need for a model adjoint, which can be taxing to develop and maintain, has increased the popularity of adjoint-free methods, in particular algorithms that use ensembles to encode flow-dependent information. The most common ensemble-based algorithms currently employed at weather centers are the ensemble Kalman filter (EnKF) (Evensen, 1994; Lorenc, 2003; Houtekamer and Zhang, 2016) and ensemble–variational assimilation (EnVar) (Buehner et al., 2010; Desroziers et al., 2014; Fairbairn et al., 2014).
@@
The assimilation of atmospheric chemical observations has been an important component of air quality modeling (Miyazaki et al., 2012b; Robichaud, 2017; Huijnen et al., 2020; Park et al., 2021; Garrigues et al., 2022), stratospheric ozone and UV forecasting (Lahoz et al., 2007; Inness et al., 2013; Rochon et al., 2019), climate modeling (Peters et al., 2010; Miyazaki et al., 2011; Bruhwiler et al., 2014; Agustí-Panareda et al., 2023), and emissions inversions (Miyazaki et al., 2012b, a; Zhu et al., 2013; Cao et al., 2020). The computational challenges associated with ensemble generation are compounded when applied to atmospheric chemistry models, which typically require more computational resources than NWP models. While many operational NWP ensemble system are run at weather centers with ensemble sizes on the order of tens to hundreds, the higher computational requirements of running atmospheric chemistry models make running a similarly designed chemical data assimilation system difficult, especially in an operational setting. When ensemble-based chemical data assimilation is feasible to run, it typically has a much smaller ensemble size than in NWP systems, often with less than a hundred members (Pendergrass et al., 2023). In contrast to most modern NWP systems, it is not uncommon for chemical data assimilation systems to use static background error covariances. For example, at Environment and Climate Change Canada (ECCC), the two operational systems that produce analyses for chemical fields, the air quality surface analysis (Ménard and Deshaies-Jacques, 2018a, b, c) and the stratospheric ozone forecasting system (Rochon et al., 2019; Buehner et al., 2024), both use static background error covariances. ECCC’s air quality forecasting system uses the Global Environmental Multiscale – Modelling Air quality and Chemistry (GEM-MACH) chemical weather model, which takes approximately four to five times more computing time and up to nine times more disk space than the Global Environmental Multiscale (GEM) NWP model that GEM-MACH is based on, illustrating the challenges with using ensemble methods with atmospheric chemical models.
@@
For our application, the ensemble of emission perturbations is computationally inexpensive to create, but translating the ensemble of emissions perturbations into an ensemble of concentration perturbations using a physically-based atmospheric model is usually very computationally demanding. EnsAI uses a U-Net convolutional neural network to map ammonia emission perturbations to ammonia surface concentration perturbations, using surface wind and temperature fields as auxiliary inputs to the neural network. EnsAI can generate the ensemble of concentrations without needing to run computationally expensive physically-based atmospheric models. As EnsAI has the upfront cost of generating an ensemble for training, as well as the training itself, the intent is to be able to use EnsAI for much longer time periods than the time period used for training, so that these initial costs are much smaller than the long term computational savings over the intended use period.
@@ 2 Emissions Inversions
This study uses the application to ammonia emission inversions as a testbed for developing the foundations of an AI-based ensemble generation system. As in Sitwell et al. (2022), ammonia retrievals from May to August 2016 from the satellite-borne CrIS instrument are used in emission inversions to refine the ammonia emissions used with the GEM-MACH air quality model. While the inversion methods are described in full in Sitwell et al. (2022), we summarize the main features in this section, as well as a few modifications made for this study.
@@ 2.1 The GEM-MACH Air Quality Model
Monthly-mean emissions, originating from either an emissions inventory or an emissions inversion, are translated into hourly emissions that can be used in GEM-MACH through the use of sector-dependent time profiles. The unperturbed set of emissions was version 3.1.2 of the operational GEM-MACH emissions data set. Figure 1 shows this emissions inventory for ammonia averaged over May to August, which are the months of interest for this study.
Figure 1: Unperturbed ammonia emissions inventory averaged over May to August. Emissions are only displayed within the domain of the EnsAI model. Particular locations examined in this work are also displayed, which are located in (in order of descending emissions) (a) North Carolina, (b) Iowa, (c) Ohio, (d) Pennsylvania, and (e) Georgia.
@@ 2.2 CrIS Ammonia Retrievals
The CrIS instrument, which is aboard the Suomi National Polar-orbiting Partnership (SNPP), NOAA-20 and NOAA-21 satellites, is a Fourier transform spectrometer that measures the infrared spectrum in three different spectral bands. The CrIS Fast Physical Retrieval (CFPR) algorithm (Shephard and Cady-Pereira, 2015; Shephard et al., 2020), based on the TES ammonia retrieval algorithm (Shephard et al., 2011), used for the ammonia retrievals makes use of the longwave band between 960 and 967 $\mathrm{cm}^{-1}$. As in Sitwell et al. (2022), we will focus on the spring and summer of 2016 (which comprises most of the peak ammonia emissions season) and use retrievals from the CrIS instrument aboard the SNPP satellite. The SNPP satellite, which is in a sun-synchronous orbit, has local overpass times at approximately 01:30 and 13:30, though only the day-time retrieval is used in the inversions. CrIS has a resolution at nadir of around 14 km. The retrievals are made on 14 pressure levels, with the averaging kernel typically peaking at or near the surface.
@@ 2.3 Inversion Methods
where $\textbf{B}_{cc}$ and $\textbf{B}_{ee}$ represent the univariate covariances for the surface concentrations and emissions, respectively, while $\textbf{B}_{ce}$ and $\textbf{B}_{ec}$ represent their cross-covariance. Although the uncertainties in the atmospheric concentration of ammonia is likely due to a number of different factors including uncertainties in emissions, meteorology, initial conditions, and chemistry, for these inversions only the uncertainty due to the emissions is considered.
@@
where $\delta\textbf{x}_{i}$ is model state perturbation of ensemble member $i$. Physically-based atmospheric models are usually computationally costly to run, so ensembles typically consist of tens to hundreds of members. These small ensemble sizes produce error covariances with low rank and spurious long-distance correlations. A localization scheme is usually included to mitigate these effects, which we implement through an element-wise multiplication (denoted by $\circ$) of the ensemble error covariance with a localization matrix $\mathcal{L}$.
@@ 3 AI Ensemble Generation
For application to ammonia emission inversions, the same perturbation to the time-mean ammonia emissions can results in different perturbations to the atmospheric ammonia concentration at different times due to the dependency on the meteorological conditions. The strength of the ammonia emissions are strongly dependent on temperature, while the local spatial distribution of its atmospheric concentration is dependent on advection patterns. This is illustrated in Figure 2, which displays the ammonia emission and surface concentration perturbations for one particular member of the ensemble. The monthly-mean ammonia emission perturbations for May is shown in the top row, while the the bottom row shows the surface concentration perturbations resulting from this emission perturbation at three different times (all at the same time, 18 UTC, but for different days). While all three surface concentration perturbations originate from the same emissions perturbation, the different meteorological conditions at each time have a clear effect on both the magnitude and spatial pattern of surface concentration perturbations.
Figure 2: Monthly-mean ammonia emission perturbations for May (top row) and the resulting ammonia surface concentration perturbations at three different time in May (bottom row) for one member of the GEM-MACH ensemble.
@@
It is worth noting that generating the ensemble of emissions is simple, as they are constructed from a Gaussian random field as described in Section 2.4. In Eq. (4), the GEM-MACH model $\mathcal{M}^{\rm{GM}}$ translates the emission ensemble into a atmospheric concentration ensemble, and the desire is for $\mathcal{M}^{\rm{EnsAI}}$ to emulate this process. This is in contrast to other AI-based geophysical systems, such as those using diffusion-based generative models (Price et al., 2023; Li et al., 2024a), that construct a latent space that is easy to sample in. In diffusion-based generative models, a diffusion process is used to add noise to an image (which in our application would correspond to an atmospheric field) until the image is dominated by noise. The goal of the noising process is for the distribution of noised images to be well described by a probability distribution that is easy to sample from. In these systems, random samples are drawn in the latent space and are then denoised using the reverse diffusion process (Song et al., 2020; Song and Ermon, 2020; Ho et al., 2020) to transform the random samples into the space of the atmospheric model. In our application, sampling from the emissions distribution is already relatively easy, so constructing and transforming into a latent space is not necessary. In both cases, samples are randomly drawn in a space where the probability distribution is relatively simple and the AI model transforms that sample to the state space of the atmospheric model. However, EnsAI can potentially be simpler than diffusion-based (or other similar) generative models since a diffusion process to denoise an image does not need to be modeled.
@@ 4 Neural Network Architecture
The input fields to EnsAI are displayed in the upper-left corner of Fig. 3, which consists of the four input fields $[\delta e,u,v,T]$ on the $256\times 256$ horizontal grid. In the encoding branch of the network, the input image is sequentially transformed into feature maps with more channels and a reduction in spatial resolution to encode different image features. The decoding branch decreases the number of channels and increases the spatial resolution, with the final output having the original spatial resolution and a single channel for the surface atmospheric concentration perturbation $\delta c$ (shown in the upper-right corner of Fig. 3). Details on each component of the U-Net can be found in Appendix B.
@@ 5 Training the Network
As the existing GEM-MACH ammonia ensemble is used for both training and testing, it must be divided into training, validation, and testing data sets. When multiple years of data are available, it is common to withhold the last year of data as a test set. However, in our case only a single spring/summer period is available that spans from May to August 2016.
Following ECCC’s operational air quality system (Ménard and Deshaies-Jacques, 2018c; Ménard et al., 2020), which evaluate model error covariances within two-month time periods, we similarly segment our ensemble into two-month durations: One period for May/June and another for July/August. We withhold the first week of data from June and August, which fall approximately at the mid-point of each two-month time period, to use as our test set. The selected test periods fall approximately after the first and third quarters of the full data set, which should be somewhat more representative of the whole period as compared to selecting the time period at the end of the data set that may be less representative of the spring/summer seasons.
Emissions inversions for ammonia are typically designed so that one inversion is done for each month. Choosing our test set to span only single weeks means that we will only be able to perform inversions using a single week’s worth of data as part of our evaluation. However, since our goal is for the EnsAI ensemble to follow the GEM-MACH ensemble as closely as possible, comparing the inversion results from the two different ensembles should be sufficient for our purpose.
A 60 member ensemble of hourly GEM-MACH model output for May to August 2016 was used for training, which amounted to $123\;\rm{days}\times 24\;\rm{h}\;\rm{day}^{-1}\times 60\;\rm{members}=177,120\;\rm{samples}$ over the whole data set. After removing the test set (about 11% of the total data), an additional 11 days (or about 9% of the total data) was reserved for the validation set used during training. The mean squared error (MSE) was used as the loss function for training. While the MSE loss function can lead to smoothing at smaller scales, no significant smoothing was observed in our experiments. Training was done on a NVIDIA A100 GPU and took approximately three days to train the network through 50 epochs. More details of the training can be found in Section S1 of the Supplement.
@@ 6 Run Time Comparison
One of the main motivations of this work is to develop a method for ensemble generation that can run significantly faster than traditional physically-based systems. The original GEM-MACH ensemble took roughly 6.5 hours of wall time to generate a week’s worth of simulated data per ensemble member when run on a CPU cluster using 720 computing nodes, with each node having access to 80 cores. In contrast, EnsAI when run on a single GPU took only 7 seconds of wall time to generate a week’s worth of simulated data per ensemble member, or about 3,300 times faster than GEM-MACH.
@@
It is important to note that while EnsAI runs significantly faster than GEM-MACH, GEM-MACH produces forecasts for hundreds of chemical fields on 85 vertical levels, while EnsAI only produces a single surface concentration field and does so on a smaller horizontal grid. However, while the wealth of information produced by GEM-MACH can have many uses, the full volume of this information is not always needed for every application. Generating an ensemble of ammonia concentrations at or near the surface to be used for emissions inversions is one example where not having as much information generated is not necessarily a draw back, but where the long run times of the physically-based model may limit how often these inversions can be performed.

### v2

@@ Abstract
Ensemble-based methods for data assimilation and emissions inversions are a popular way to encode flow-dependency within the model error covariance. While most ensemble methods do not require the use of an adjoint model, the need to repeatedly run a geophysical model to generate the ensemble can be a significant computational burden. In this paper, we introduce EnsAI, a new AI-based ensemble generation system for atmospheric chemical constituents. When trained on an existing ensemble for ammonia generated by the GEM-MACH air quality model, it was shown that the ensembles produced by EnsAI can accurately reproduce the meteorology-dependent features of the original ensemble, while generating the ensemble 3,300 times faster than the original GEM-MACH ensemble. While EnsAI requires an upfront cost for generating the ensemble used for training, as well as the training itself, the long term computational savings can greatly exceed these initial computational costs. When used in an emissions inversion system, EnsAI produced similar inversion results to those in which the original GEM-MACH ensemble was used while using significantly less computational resources.
@@ 1 Introduction
While having an accurate estimate of a model’s error covariance is of high importance in many applications, actually obtaining such an estimate is often difficult. Earlier numerical weather prediction (NWP) assimilation systems often used static forecast error covariances as this was more computationally feasible at the time. But as more computational resources became available, more weather centers were able to incorporate flow-dependent error covariances within their systems. One popular method for doing this was by using four-dimensional variational (4DVar) assimilation (Thepaut and Courtier, 1991; Zupanski, 1993; Rabier et al., 2000). However, 4DVar’s need for a model adjoint, which can be taxing to develop and maintain, has increased the popularity of adjoint-free methods, in particular algorithms that use ensembles to encode flow-dependent information. The most common ensemble-based algorithms currently employed at weather centers are the ensemble Kalman filter (EnKF) (Evensen, 1994; Lorenc, 2003; Houtekamer and Zhang, 2016) and ensemble–variational assimilation (EnVar) (Buehner et al., 2010; Desroziers et al., 2014; Fairbairn et al., 2014).
@@
The assimilation of atmospheric chemical observations has been an important component of air quality modeling (Miyazaki et al., 2012b; Robichaud, 2017; Huijnen et al., 2020; Park et al., 2021; Garrigues et al., 2022), stratospheric ozone and UV forecasting (Lahoz et al., 2007; Inness et al., 2013; Rochon et al., 2019), climate modeling (Peters et al., 2010; Miyazaki et al., 2011; Bruhwiler et al., 2014; Agustí-Panareda et al., 2023), and emissions inversions (Miyazaki et al., 2012b, a; Zhu et al., 2013; Cao et al., 2020). The computational challenges associated with ensemble generation are compounded when applied to atmospheric chemistry models, which typically require more computational resources than NWP models. While many operational NWP ensemble systems are run at weather centers with ensemble sizes on the order of tens to hundreds, the higher computational requirements of running atmospheric chemistry models make running a similarly designed chemical data assimilation system difficult, especially in an operational setting. When ensemble-based chemical data assimilation is feasible to run, it typically has a much smaller ensemble size than in NWP systems, often with less than a hundred members (Pendergrass et al., 2023). In contrast to most modern NWP systems, it is not uncommon for chemical data assimilation systems to use static background error covariances. For example, at Environment and Climate Change Canada (ECCC), the two operational systems that produce analyses for chemical fields, the air quality surface analysis (Ménard and Deshaies-Jacques, 2018a, b, c) and the stratospheric ozone forecasting system (Rochon et al., 2019; Buehner et al., 2024), both use static background error covariances. ECCC’s air quality forecasting system uses the Global Environmental Multiscale – Modelling Air quality and Chemistry (GEM-MACH) chemical weather model, which takes approximately four to five times more computing time and up to nine times more disk space than the Global Environmental Multiscale (GEM) NWP model that GEM-MACH is based on, illustrating the challenges with using ensemble methods with atmospheric chemical models.
@@
For our application, the ensemble of emissions perturbations is computationally inexpensive to create, but translating the ensemble of emissions perturbations into an ensemble of atmospheric concentration perturbations using a physically-based atmospheric model is usually very computationally demanding. EnsAI uses a U-Net convolutional neural network to map ammonia emissions perturbations to ammonia surface concentration perturbations, using surface wind and temperature fields as auxiliary inputs to the neural network. EnsAI can generate the ensemble of concentrations without needing to run computationally expensive physically-based atmospheric models. As EnsAI has the upfront cost of generating an ensemble for training, as well as the training itself, the intent is to be able to use EnsAI for much longer time periods than the time period used for training, so that these initial costs are much smaller than the long term computational savings over the intended use period.
@@ 2 Emissions Inversions
This study uses the application to ammonia emissions inversions as a testbed for developing the foundations of an AI-based ensemble generation system. As in Sitwell et al. (2022), ammonia retrievals from the satellite-borne CrIS instrument are used in emissions inversions to refine the ammonia emissions used with the GEM-MACH air quality model. While the inversion methods are described in full in Sitwell et al. (2022), we summarize the main features in this section, as well as a few modifications made for this study.
@@ 2.1 The GEM-MACH Air Quality Model
Hourly ammonia emissions are derived by combining monthly-mean ammonia emissions with meteorological information in the bidirectional flux model. The emissions for all other species are computed by translating monthly or annual emissions inventories into hourly emissions through the use of sector-dependent time profiles. The unperturbed set of emissions was version 3.1.2 of the operational GEM-MACH emissions data set. The annual mean of this emissions inventory for ammonia in shown in Figure 1.
Figure 1: Annual mean ammonia emissions for the (unperturbed) inventory. Emissions are only displayed within the domain of the EnsAI model. Particular locations examined in this work are also displayed, which are located in (in order of descending emissions) (a) North Carolina, (b) Iowa, (c) Ohio, (d) Pennsylvania, and (e) Georgia.
@@ 2.2 CrIS Ammonia Retrievals
The CrIS instrument, which is aboard the Suomi National Polar-orbiting Partnership (SNPP), NOAA-20 and NOAA-21 satellites, is a Fourier transform spectrometer that measures the infrared spectrum in three different spectral bands. The CrIS Fast Physical Retrieval (CFPR) algorithm (Shephard and Cady-Pereira, 2015; Shephard et al., 2020), based on the TES ammonia retrieval algorithm (Shephard et al., 2011), used for the ammonia retrievals makes use of the longwave band between 960 and 967 $\mathrm{cm}^{-1}$. As we will focus on the years 2014 to 2016, we use retrievals from the CrIS instrument aboard the SNPP satellite. The SNPP satellite, which is in a sun-synchronous orbit, has local overpass times at approximately 01:30 and 13:30, though only the day-time retrieval is used in the inversions. CrIS has a resolution at nadir of around 14 km. The retrievals are made on 14 pressure levels, with the averaging kernel typically peaking at or near the surface.
@@ 2.3 Inversion Methods
where $\textbf{B}_{cc}$ and $\textbf{B}_{ee}$ represent the univariate covariances for the surface concentrations and emissions, respectively, while $\textbf{B}_{ec}$ represents their cross-covariance. Although the uncertainties in the atmospheric concentration of ammonia is likely due to a number of different factors including uncertainties in emissions, meteorology, initial conditions, and chemistry, for these inversions only the uncertainty due to the emissions is considered.
@@
where $\delta\textbf{x}_{i}$ is the model state perturbation of ensemble member $i$. Physically-based atmospheric models are usually computationally costly to run, so ensembles typically consist of tens to hundreds of members. These small ensemble sizes produce error covariances with low rank and spurious long-distance correlations. A localization scheme is usually included to mitigate these effects, which we implement through an element-wise multiplication (denoted by $\circ$) of the ensemble error covariance with a localization matrix $\mathcal{L}$.
@@ 3 AI Ensemble Generation
For application to ammonia emissions inversions, the same perturbation to the time-mean ammonia emissions can results in different perturbations to the atmospheric ammonia concentration at different times due to the dependency on the meteorological conditions. The strength of the ammonia emissions are strongly dependent on temperature, while the local spatial distribution of its atmospheric concentration is dependent on advection patterns. This is illustrated in Figure 2, which displays the ammonia emission and surface concentration perturbations for one particular member of the ensemble. The monthly-mean ammonia emission perturbations for May is shown in the top row, while the the bottom row shows the surface concentration perturbations resulting from this emission perturbation at three different times (all at the same time, 18 UTC, but for different days). While all three surface concentration perturbations originate from the same emissions perturbation, the different meteorological conditions at each time have a clear effect on both the magnitude and spatial pattern of surface concentration perturbations.
Figure 2: Monthly-mean ammonia emissions perturbations for May (top row) and the resulting ammonia surface concentration perturbations at three different time in May (bottom row) for one member of the GEM-MACH ensemble.
@@
It is worth noting that generating the ensemble of emissions is simple, as they are constructed from a Gaussian random field as described in Section 2.4. In Eq. (4), the GEM-MACH model $\mathcal{M}^{\rm{GM}}$ translates the emissions ensemble into an atmospheric concentration ensemble, and the desire is for $\mathcal{M}^{\rm{EnsAI}}$ to emulate this process. This is in contrast to other AI-based geophysical systems, such as those using diffusion-based generative models (Price et al., 2023; Li et al., 2024a), that construct a latent space that is easy to sample in. In diffusion-based generative models, a diffusion process is used to add noise to an image (which in our application would correspond to an atmospheric field) until the image is dominated by noise. The goal of the noising process is for the distribution of noised images to be well described by a probability distribution that is easy to sample from. In these systems, random samples are drawn in the latent space and are then denoised using the reverse diffusion process (Song et al., 2020; Song and Ermon, 2020; Ho et al., 2020) to transform the random samples into the space of the atmospheric model. In our application, sampling from the emissions distribution is already relatively easy, so constructing and transforming into a latent space is not necessary. In both cases, samples are randomly drawn in a space where the probability distribution is relatively simple and the AI model transforms that sample to the state space of the atmospheric model. However, EnsAI can potentially be simpler than diffusion-based (or other similar) generative models since a diffusion process to denoise an image does not need to be modeled.
@@ 4 Neural Network Architecture
The input fields to EnsAI are displayed in the upper-left corner of Fig. 3, which consists of the four input fields $\delta e,u,v,T]$ on the $256\times 256$ horizontal grid. In the encoding branch of the network, the input image is sequentially transformed into feature maps with more channels and a reduction in spatial resolution to encode different image features. The decoding branch decreases the number of channels and increases the spatial resolution, with the final output having the original spatial resolution and a single channel for the surface atmospheric concentration perturbation $\delta c$ (shown in the upper-right corner of Fig. 3). Details on each component of the U-Net can be found in Appendix B.
@@ 5 Training the Network
An existing GEM-MACH ammonia ensemble for January to December 2016 was used for training the neural network. The ensemble consisted of 60 members of hourly GEM-MACH model output, which amounted to $366\;\rm{days}\times 24\;\rm{h}\;\rm{day}^{-1}\times 60\;\rm{members}=527,040\;\rm{samples}$ for the training set. The mean squared error (MSE) was used as the loss function for training. While the MSE loss function can lead to smoothing at smaller scales, no significant smoothing was observed in our experiments. Training was done on a NVIDIA A100 GPU and took approximately 18 days to train the network through 50 epochs. More details of the training can be found in Section S1 of the Supplement.
Ensembles for January to December 2015 and July 2014 were generated to use as test and validation data sets, respectively (although we note that the test set covers slightly less than a full year’s worth of data, as the data for January 9 to 21 were excluded from this set due to unavailable GEM-MACH files during these times).
@@ 6 Run Time Comparison
One of the main motivations of this work is to develop a method for ensemble generation that can run significantly faster than traditional physically-based systems. The original GEM-MACH ensemble took roughly 6.5 hours of wall time to generate a week’s worth of simulated data per ensemble member when run on a CPU cluster using 720 computing nodes. In contrast, EnsAI when run on a single GPU took only 7 seconds of wall time to generate a week’s worth of simulated data per ensemble member, or about 3,300 times faster than GEM-MACH.
@@
It is important to note that while EnsAI runs significantly faster than GEM-MACH, GEM-MACH produces forecasts for hundreds of chemical fields on 85 vertical levels, while EnsAI only produces a single surface concentration field. However, while the wealth of information produced by GEM-MACH can have many uses, the full volume of this information is not always needed for every application. Generating an ensemble of ammonia concentrations at or near the surface to be used for emissions inversions is one example where not having as much information generated is not necessarily a draw back, but where the long run times of the physically-based model may limit how often these inversions can be performed.